# Enabling Quick, Accurate Crowdsourced Annotation for Elevation-Aware Flood Extent Mapping

### v1

@@ Abstract
In order to assess damage and properly allocate relief efforts, mapping the extent of flood events is a necessary and important aspect of disaster management. In recent years, deep learning methods have evolved as an effective tool to quickly label high-resolution imagery and provide necessary flood extent mappings. These methods, though, require large amounts of annotated training data to create models that are accurate and robust to new flooded imagery. In this work, we provide FloodTrace, an application that enables effective crowdsourcing for flooded region annotation for machine learning training data, removing the requirement for annotation to be done solely by researchers. We accomplish this through two orthogonal methods within our application, informed by requirements from domain experts. First, we utilize elevation-guided annotation tools and 3D rendering to inform user annotation decisions with digital elevation model (DEM) data, improving users’ annotation accuracy. For this purpose, we provide a unique annotation method that uses topological data analysis to outperform the state-of-the-art elevation-guided annotation tool in efficiency. Second, we provide a framework for researchers to review aggregated crowdsourced annotations and correct inaccuracies using methods inspired by uncertainty visualization. We conducted a user study to confirm our application’s effectiveness in which 266 graduate students annotated high-resolution aerial imagery from Hurricane Matthew in North Carolina. Experimental results show the accuracy and efficiency benefits of our application apply even for untrained users. In addition, using our aggregation and correction framework, flood detection models trained on crowdsourced annotations were able to achieve performance equal to models trained on expert-labeled annotations, while requiring a fraction of the time on the part of the researcher. Our solution is completely web-based for ease of access with a demo available at https://ldyken53.github.io/threejs-floodmap/.
@@
An annotation made on a 6500 $\times$ 3000 pixel aerial image of flooding caused by Hurricane Matthew using only ten clicks of our topological segmentation tool. Pixels overlaid with blue have been labeled as dry and red as flooded. Note that areas labeled as flooded can appear green on the left, this is due to the tree canopy obscuring the flooded ground beneath. This tool guides annotation using the corresponding elevation data for a region, resulting in accurate labels. This can be seen by the highlighted areas where the tool annotated flooding around small dry features, but left them unlabeled. In our experimental user study, we show that this tool is more efficient than the state-of-the-art elevation-guided tool and can be used effectively by untrained participants.
@@ Introduction
Existing flood annotation datasets are most commonly created by manual labeling of aerial imagery with simple brushes and polygon selection tools, which is extremely time-consuming. Semi-automatic approaches improve a user’s productivity by allowing them to label larger areas more quickly with the assistance of a guiding algorithm. Semi-automatic tools have been used previously to create flood annotation datasets [24, 8, 51, 42] through various algorithms on the flooded imagery being labeled. While these tools can improve productivity, our work builds on insight from domain experts which motivates a focus on elevation-guided annotation of flooded regions. The state-of-the-art machine learning models in this domain are physics-guided and take into account the elevation of the region being observed by utilizing Digital Elevation Model (DEM) data [68, 37, 34, 46, 53]. These models require very accurate ground truth labels for training. This is especially true for areas of high and low relative elevation in a dataset, and for adjacent pixels with very different elevation, named border pixels. With these models, a high-elevation pixel that is mislabeled as flooded can cause many surrounding pixels to be incorrectly inferred as flooded. By supplementing aerial imagery data with corresponding DEM data, the annotation process can be guided by elevation and lead to more accurate training data for dependent models. Moreover, elevation data can inform the labeling of otherwise ambiguous pixels covered by clouds and tree canopies.
@@
In this work, we propose FloodTrace, a system for quick, accurate annotation that improves researchers’ workflows and enables high-quality crowdsourcing for flood extent mapping. Our design was informed by insight from domain experts in the form of repeated meetings and demos with one senior and two junior researchers working on machine learning models for flood extent mapping. Because flooding is naturally related to the varying elevation of land in an area, we take advantage of DEM data to inform annotation decisions with 3D views and tools, building a fully elevation-guided annotation application. We provide state-of-the-art and novel semi-automatic annotation tools that utilize elevation data and topological data analysis to increase annotation speed and accuracy. In our evaluation, we show that these tools can be used effectively by untrained students as proof of crowdsourcing potential.
@@
FloodTrace is designed to encourage and improve the quality of crowdsourcing for this domain. Because of this, our solution is implemented as an interactive web application, making it easily accessible for end users. To improve researcher trust and understanding, our application provides an uncertainty visualization component to visualize aggregated crowdsourced annotations and find areas of high uncertainty between annotators. Annotation tools can then be used to directly correct those regions, improving the quality of aggregated annotations as training data.
@@
To evaluate FloodTrace, we conducted a user study with 266 computer science graduate students in which we logged all interactions and results over several annotation tasks. Aggregated annotations were then used to train elevation-aware machine learning models, which were tested for flood detection on unseen regions to assess training data quality. We find that elevation-guided tools increased the accuracy of participants’ annotations, while our topological segmentation tool greatly increased annotation efficiency. We show that by using our aggregate visualization and correction workflow, crowdsourced annotations can be used to create models with equal performance to models trained on expert-labeled annotations, while requiring a fraction of the effort on the part of the researcher. Our contributions are:
  • A novel framework for accurately annotating flooded regions using both satellite imagery and elevation data. Effective tools in a web-based application enable easy and productive crowdsourcing.
  • A visualization component for reviewing and correcting crowdsourced annotations by showing uncertainty between annotators.
  • An evaluation of our framework’s performance through a user study with 266 participants.
  • The dataset of annotations from the user study, consisting of 1,321 unique annotations on eight distinct regions of North Carolina during Hurricane Matthew. Average dimensions of each annotation were $2573\times 1284$ pixels. Annotations are provided with their metadata at [anonymized link].
@@ 1 Background and Related Work
In this section, we first discuss the workflow of machine learning researchers to provide context for how our application can improve existing methods for flood extent mapping (subsection 1.1). Then, we investigate previous work in crowdsourcing applications to motivate and inform our design (subsection 1.2) and review related work in annotation tools (subsection 1.3). Finally, we motivate and give background for the topological data analysis methods used for our novel elevation-guided annotation tool (subsection 1.4).
@@ 1.1 ML for Flood Extent Mapping Workflow
While workflows can vary based on the specific flooding application of the models being trained, we report here an example workflow of training ML models for flood extent mapping from a discussion with our expert collaborators.
@@
The process begins with data collection, acquiring aerial imagery taken during a flood disaster with corresponding digital elevation model (DEM) data. This imagery can be obtained from a source such as NOAA’s National Geodetic Survey of Emergency Response Imagery [47] in the form of patches with global Coordinate Reference System (CRS) info. These patches are then stitched together into test regions by loading them into QGIS[50], finding which patches are needed to create a test region using their CRS info, and then using the GDAL[23] library in Python to combine them into one image. DEM data can then be created for that region in the same way by downloading and stitching together data patches at corresponding CRS values, with DEM data widely available for locations in the US[25] and globally[1]. Afterward, both imagery and DEM data need to be resampled into the same spatial resolution so that the pixels are perfectly aligned.
@@
The next step is to annotate the flooded and dry areas in the imagery to obtain ground truth labels for dependent model training and performance evaluation. Study areas for flood mapping can easily have millions of pixels, so while this can be done with semi-automatic or manual tools, it is still the most time-consuming part of the flood extent mapping process, taking multiple hours per region. Once the annotation is complete, the labeled data can be used to train deep learning models for flood detection tasks such as segmenting new data into flooded and dry regions. In order to create ML models that can robustly detect flooding on unseen data, large training sets are required. While existing work has explored using fully automated methods for creating labeled flood datasets[7], they are not widely used in practice because of quality concerns and the added difficulty for researchers to understand and explain model behavior. In our collaborators’ lab, the bulk of annotation is accomplished by trained graduate students. In our work, we hope to improve the workflows of ML researchers by reducing the burden of annotation through efficient semi-automatic tools and enabling crowdsourcing.
@@ 1.2 Crowdsourcing for Research
Nowadays, many projects (e.g. eBird[66], GalaxyZoo[43]) have shown that crowdsourcing and citizen science can be successful for large-scale data collection and annotation. Because any barrier to entry will lower the number of users of a platform, crowdsourcing applications are almost always web-based for ease of access. To define the qualities of successful crowdsourcing projects, Law et al. [41] deduced a set of requirements that make crowdsourcing feasible, desirable, and useful for a given research problem. Flood extent mapping satisfies their challenges of data sensitivity, quantity, and availability, along with crowd interest, intention, and ability as proven by other applications around crowdsourced flood event data[58, 48, 57]. In addition, previous work[42] has shown that, while individual crowdworkers may not perform well in flood mapping, aggregating annotations can lead to high-quality results. This work aggregates user annotations by labeling patches as flooded or dry if the ratio of received labels for that patch is above an empirically chosen certainty threshold. FloodTrace is unique in providing a system for interactively visualizing aggregated annotations with the uncertainty between annotators, which allows experts to quickly revise and improve the set of annotations.
@@ 1.3 Annotation Tools
Multiple web-based libraries such as OpenStreetMap[27], QGIS Cloud[50], and ArcGIS Online[3] provide tools for geospatial data. Each of these tools was made for different use cases but can support DEM data, interactive 3D rendering, and basic annotation with text, polygons, and brushes. Their potential for flood extent mapping, however, is limited by the fact that none provide semi-automatic annotation tools for the purpose. Even with crowdsourcing, manual annotation of flooded imagery with polygon selections and brushes is too time-consuming to be desirable and results in less accurate training data.
@@
For semi-automatic annotation, there are two main classes of tools in this domain. The first are those that produce a segmentation of the input and allow users to apply flood or dry class labels to selected segments (segmentation-type tools). Liang et al. [42] use imagery patches exported from a graph-based clustering approach as these segments, and show that these patches can be labeled effectively by non-experts. Other work has used image features produced by class-agnostic neural networks for labeling to create large flood annotation datasets [8, 51]. The second class of tools are those where the user selects seed pixels which are then extended with connected pixels by some rule (extension-type tools). As an example, Gebrehiwot et al. [24] use a tool to automatically label connected pixels of the same color as flooded or dry. While the semi-automatic annotation tools mentioned so far operate directly on aerial imagery data, recent exploratory work has been done on creating an elevation-guided extension-type tool for labeling flooded datasets [2]. This work utilizes DEM data to select all connected downstream (lower elevation) or upstream (higher elevation) pixels from a selected seed pixel, which they call the breadth-first search (BFS) method. As our design seeks to enable elevation-guided annotation, we adapt their BFS method as a state-of-the-art tool in our application (subsubsection 3.2.1). Still, it is important to note that the original work[2] does not support interactive 3D rendering, can only operate on small data patches, and has poor performance, all issues that are addressed by our web-based solution with GPU accelerated rendering. In addition, we develop a new elevation-guided tool that uses topological data analysis to annotate regions much more quickly than the BFS tool.
Figure 1: Example of contour tree and Morse-Smale complex (MSC) segmentation on a $250\times 250$ slice of DEM elevation data using the Topology Toolkit in Paraview. Before creating either segmentation, simplification by persistence was applied with a threshold of 0.1 of the function range. The contour tree segmentation shows its clear use case as a tool for selecting terrain features such as the hills and river in this example.
1.4 Topological Analysis of DEM Data
@@
As the size and resolution of DEM data grow, the need for generalization and simplification of features is well known; multi-scale views are required in order to conduct meaningful analysis on DEM landforms [64, 65, 67, 31, 18]. Several methods have been proposed for this, which needed generalization and simplification of DEM data. Yu et al. [65] propose a method combining structural analysis and statistical filtering for simplification while preserving smooth morphology and structural edges. Guilbert [31] uses contours to generate a feature tree for underlying data at multiple levels of detail. Wu et al. [67] use a localized contour tree method to detect and characterize surface depressions across scales. Recently, Corcoran et al. [9] show the advantages of persistent homology methods in terrain analysis for their robustness to noise and facilitating of machine learning methods utilizing persistence diagrams as signatures of a dataset. In our work, we leverage persistent homology and contour tree methods on DEM data to enable robust annotation of features at different levels of detail.
By extracting the underlying structural information from scalar data, topological data analysis (TDA) provides many opportunities for data segmentation and simplification. Topological methods have been successfully used to extract features of interest in a variety of scientific domains [5, 6, 40, 59, 49, 32, 33, 52]. These data features can be effectively identified by using topological structures such as the Morse-Smale complex (MSC) [16] and contour trees to create segmentations of the data. The MSC does this through gradient flow, segmenting the scalar field data into regions consisting of all the paths connecting a local maximum and local minimum that follow the gradient of the scalar field. Contour trees segment data by considering contours through a sweep of the function range, defining segments as the birth or split of a contour until it merges with another or vanishes. We provide an example of both methods of TDA segmentation on a small piece of DEM data in Figure 1. As shown, our use case of selecting dry and flooded features correlates well with the contour tree segmentation. This method identifies terrain features such as hills and rivers, along with contours at various elevations which allow for the selection of the precise water level of a flooded area. In this work, we propose using the contour tree to create a novel segmentation-type tool for elevation-guided annotation for flood extent mapping.
@@
One challenge is that, because DEM data is complex and large-scale, contour tree segmentations on raw data are mostly made of noisy and small segments. This, along with the necessity of identifying features at varying levels of detail, motivates our use of persistent homology for data simplification. Persistent homology [17] is a powerful tool that can be used to remove topological features that are unimportant or correspond to low-intensity noise. This is done by considering the persistence of these features, including their lifespan between birth and death when conducting a sweep through the range of the function. Simplification by persistence involves removing critical points corresponding to the birth and death of a topological feature if that feature’s lifespan is below some threshold. In doing this, simplification by persistence preserves salient features in the data, in contrast to other simplification methods like downsampling the spatial resolution. As an example, a spatially small feature with a huge variance in elevation compared to its surroundings would have a large persistence value, and not be simplified until high thresholds, whereas downsampling the data’s spatial resolution would remove this feature. In this way, our simplification is more robust for maintaining topological features in high-resolution data. By simplifying our data at varying persistence thresholds, we can provide effective multiscale data for selecting features in our application; this affords users the ability to quickly annotate large, obviously flooded and dry features, while maintaining the detail necessary to accurately label smaller features.
@@
Similar topological techniques have previously been utilized for semi-automatic annotation in neuron tracing of large brain volumetric data [45]. This work utilizes the MSC to extract ridge-like structures which correspond to neuron center-lines, which can then guide user tracing. Simplification by persistence is used with an empirically chosen persistence threshold to remove noise and create a sparser dataset for users to interact with. While taking inspiration from this work, our system adapts topological methods specifically to the labeling of flood extent in terrain data. Our method of selecting contour tree segmentations that correspond to features below or above water level in terrain data is unique. Furthermore, our method of allowing interactive selection of multiscale features by providing data simplified at varying persistence thresholds is a novel technique for annotation and is necessary to find landforms at different scales within DEM data. Finally, FloodTrace is the only system the authors found that integrates topological segmentation into a web-based application for interaction.
Figure 2: An illustration of the full processing pipeline for FloodTrace.

### v2

# Enabling Fast and Accurate Crowdsourced Annotation for Elevation-Aware Flood Extent Mapping

@@ Abstract
Mapping the extent of flood events is a necessary and important aspect of disaster management. In recent years, deep learning methods have evolved as an effective tool to quickly label high-resolution imagery and provide necessary flood extent mappings. These methods, though, require large amounts of annotated training data to create models that are accurate and robust to new flooded imagery. In this work, we present FloodTrace, a web-based application that enables effective crowdsourcing of flooded region annotation for machine learning applications. To create this application, we conducted extensive interviews with domain experts to produce a set of formal requirements. Our work brings topological segmentation tools to the web and greatly improves annotation efficiency compared to the state-of-the-art. The user-friendliness of our solution allows researchers to outsource annotations to non-experts and utilize them to produce training data with equal quality to fully expert-labeled data. We conducted a user study to confirm our application’s effectiveness in which 266 graduate students annotated high-resolution aerial imagery from Hurricane Matthew in North Carolina. Experimental results show the efficiency benefits of our application for untrained users, with median annotation time less than half the state-of-the-art annotation method. In addition, using our application’s aggregation and correction framework, flood detection models trained on crowdsourced annotations were able to achieve performance equal to models trained on fully expert-labeled annotations, while requiring a fraction of the expert’s time.
keywords:
Terrain visualization, machine learning, collaborative and distributed visualization, topology-based techniques
@@
[…]
@@ Introduction
In this domain, annotation consists of assigning categories (such as flooded or dry) to the pixels of an image, usually through directly applying color. Existing flood annotation datasets are most commonly created by manual labeling of aerial imagery with simple brushes and polygon selection tools, which is extremely time-consuming. Semi-automatic approaches improve a user’s productivity by allowing them to label larger areas more quickly with the assistance of a guiding algorithm. Semi-automatic tools have been used previously to create flood annotation datasets [26, 9, 51, 42] through various algorithms on the imagery being labeled. While these tools can improve efficiency, our work builds on insight from domain experts which motivates a focus on elevation-guided annotation of flooded regions. Recently, elevation guided methods have become more common, where satellite imagery is supplemented with a corresponding elevation model to improve the accuracy of flood extent mapping. The state-of-the-art machine learning models developed for this task are elevation-guided and use physics-informed algorithms that utilize Digital Elevation Model (DEM) data [67, 38, 35, 46, 53]. These models require very accurate ground truth labels for training. This is especially true for areas of high and low relative elevation in a dataset, and for adjacent pixels with different elevation, named border pixels. With these models, a high-elevation pixel that is mislabeled as flooded can cause many surrounding pixels to be incorrectly inferred as flooded. By using both aerial imagery and corresponding elevation data, elevation-guided annotation can create more accurate training data for dependent models.
@@
In this paper, we present FloodTrace, an interactive web-based system for quick, accurate annotation that improves researchers’ workflows and enables high-quality crowdsourced labeling for flood extent mapping. Our design was informed by domain experts through several meetings and demos with one senior and two junior researchers working on machine learning models for flood extent mapping. In our system, we take advantage of elevation data to inform annotation decisions with 3D interactive visualization and semi-automatic annotation algorithms.
@@
FloodTrace is designed to encourage and improve the quality of crowdsourcing for this domain. For this goal, our solution is implemented as an interactive web application, making it easily accessible for end users. Our systems brings for the first time topological data analysis to a convenient web environment; it provides the ability to select features based on topological segmentations, a process that would otherwise require complex setup, expertise with topological toolkits, and slow, exhaustive workflows. To improve researcher trust and understanding, our application provides an uncertainty visualization component to visualize aggregated crowdsourced annotations and find areas of high uncertainty between annotators. Our annotation tools can be used to directly correct labels for these uncertain areas, quickly improving the quality of the aggregated annotations.
@@
To evaluate FloodTrace, we conducted a user study with 266 computer science graduate students in which we logged all interactions and results over several annotation tasks. Aggregated annotations were then used to train elevation-guided machine learning models, which were tested for flood detection on unseen regions to assess training data quality. In our experimentation, we found that elevation-guided tools increased the accuracy of participants’ annotations, while our topological segmentation tool greatly increased annotation efficiency. We show that by using our aggregate visualization and correction workflow, crowdsourced annotations can be used to create models with equal performance to models trained on fully expert-labeled annotations, while requiring a fraction of an expert’s time. In summary, our contributions are:
  • An intuitive web-based tool for crowdsourced annotation of flood inundation maps that leverages satellite imagery, elevation data, and topological segmentation to enable non-experts to quickly and accurately label flooded areas. Our application is the first to bring users access to topological segmentation techniques on the web.
  • A novel review tool for visualizing aggregated crowdsourced flood annotations that enables interactive correction to improve labels.
  • An evaluation of our framework through a user study with 266 participants, producing a total of 1,321 unique annotations over eight distinct regions of North Carolina flood areas during Hurricane Matthew.
@@ 1 Background and Related Work
In this section, we first discuss the workflow of machine learning researchers to provide context for how our application can improve existing methods for flood extent mapping (subsection 1.1). Then, we explore previous work in crowdsourcing applications to motivate and inform our design (subsection 1.2) and review related work in annotation tools (subsection 1.3). Finally, we motivate and give background for the topological data analysis methods used for our novel elevation-guided annotation tool (subsection 1.4).
@@ 1.1 ML for Flood Extent Mapping Workflow
While workflows can vary based on the specific application, we report here an example workflow of training models for flood extent mapping from a discussion with our expert collaborators.
@@
The process begins with data collection, acquiring aerial imagery taken during a flood disaster with corresponding digital elevation model (DEM) data. This imagery can be obtained from a source such as NOAA’s National Geodetic Survey of Emergency Response Imagery [47] in the form of patches with global Coordinate Reference System (CRS) info. These patches are then stitched together into test regions by loading them into QGIS[50], finding which patches are needed to create a test region using their CRS info, and then using the GDAL[25] library in Python to combine them into one image. DEM data can then be created for that region in the same way by downloading and stitching together data patches at corresponding CRS values, with DEM data widely available for locations in the US[27] and globally[1]. Afterward, both imagery and DEM data need to be resampled into the same spatial resolution so that the pixels are aligned.
@@
The next step is to annotate the flooded and dry areas in the imagery to obtain ground truth labels for model training and evaluation. Study areas for flood mapping can easily have millions of pixels, so while this can be done with semi-automatic tools, it is still the most time-consuming part of the flood extent mapping process, taking multiple hours per region. In order to create ML models that can robustly detect flooding on unseen data, large training sets are required. While existing work has explored using fully automated methods for creating labeled flood datasets[7], they are not widely used in this domain because of quality concerns and the added difficulty for researchers to understand and explain model behavior after training. In our collaborators’ lab, the bulk of annotation is accomplished by trained graduate students. In our work, we hope to improve the workflows of ML researchers by reducing the burden of annotation through crowdsourcing and more efficient semi-automatic tools.
@@ 1.2 Crowdsourcing for Research
Many projects (e.g. eBird[65], GalaxyZoo[43]) have shown that crowdsourcing and citizen science can be successful for large-scale data collection and annotation. Because any barrier to entry will lower the number of users of a platform, crowdsourcing applications are almost always web-based for ease of access. To define the qualities of successful crowdsourcing projects, Law et al. [41] deduced a set of requirements that make crowdsourcing feasible, desirable, and useful for a given research problem. Flood extent mapping satisfies their challenges of data sensitivity, quantity, and availability, along with crowd interest, intention, and ability as proven by other applications around crowdsourced flood event data[58, 48, 57]. In addition, previous work[42] has shown that, while individual crowdworkers may not perform well in flood mapping, aggregating annotations can lead to high-quality results. This work aggregates user annotations by labeling patches as flooded or dry if the ratio of received labels for that patch is above an empirically chosen certainty threshold. While we experimented with this method of aggregating annotations, we found better ML model performance by training using scored soft labels of flood and dry created from averaging user annotations. Within crowdsourcing tools, FloodTrace is unique in providing a system for experts to quickly visualize areas of uncertainty between annotators and apply new labels to revise and improve the set of annotations.
@@ 1.3 Annotation Tools
Multiple libraries such as OpenStreetMap[29], QGIS Cloud[50], and ArcGIS Online[3] provide annotation tools for geospatial data. Each of these tools was made for different use cases but can support DEM data, interactive 3D rendering, and basic annotation with text, polygons, and brushes. Their potential for flood extent mapping, however, is limited by the fact that none provide semi-automatic annotation tools for the purpose. Even with crowdsourcing, manual annotation of flooded imagery with polygon selections and brushes is too time-consuming to be desirable and results in less accurate training data.
@@
For semi-automatic annotation, there are two main classes of tools in this domain. The first are those that produce a segmentation of the input and allow users to apply flood or dry class labels to selected segments (segmentation-type tools). Liang et al. [42] use imagery patches exported from a graph-based clustering approach as these segments, and show that these patches can be labeled effectively by non-experts. Other work has used image features produced by class-agnostic neural networks for labeling to create large flood annotation datasets [9, 51]. The second class of tools are those where the user selects seed pixels which are then extended with connected pixels by some rule (extension-type tools). As an example, Gebrehiwot et al. [26] use a tool to automatically label connected pixels of the same color as flooded or dry. While the semi-automatic annotation tools mentioned so far operate directly on aerial imagery data, recent exploratory work has been done on creating an elevation-guided extension-type tool for labeling flooded datasets [2]. This work utilizes DEMs to select all connected downstream (lower elevation) or upstream (higher elevation) pixels from a selected seed pixel, which they call the breadth-first search (BFS) method. We adapt their BFS method as a state-of-the-art tool in our application (subsubsection 3.2.1). It is important to note that the original work[2] does not support interactive 3D rendering, can only operate on small data patches, and requires installation, all issues that are addressed by our web-based solution with GPU accelerated rendering.
1.4 Topological Analysis of Elevation Data
@@
Topological methods provide powerful tools for analyzing elevation data features that are critical for flood mapping. Previously, they have been successfully used to simplify and extract features of interest in a variety of scientific domains [5, 6, 49, 34, 52] and specifically applied to DEM data. Yu et al. [64] propose a method combining structural analysis and statistical filtering for terrain simplification while preserving smooth morphology and structural edges. Guilbert [33] uses contours to generate a feature tree for underlying data at multiple levels of detail. Wu et al. [66] use a localized contour tree method to detect and characterize surface depressions across scales. Recently, Corcoran et al. [10] show the advantages of persistent homology methods in terrain analysis for their robustness to noise and facilitating of machine learning methods.
@@
The challenge in analyzing elevation data lies in its multi-scale nature, with terrain features existing at widely varying scales. This makes it difficult to identify relevant features, making multi-scale views required in order to conduct meaningful analysis [63, 64, 66, 33, 20]. To address this challenge, we utilize persistent homology, a mathematical framework that quantifies the importance of topological features across different scales. Persistent homology is a powerful tool to remove topological features smaller than a given scale [19]. This is done by considering the persistence of these features, which is their lifespan when conducting a sweep through the range of the function. Simplification by persistence involves removing a topological feature if its lifespan is below some threshold. In doing this, simplification by persistence is robust for maintaining topological features in high-resolution data and preserves salient details. As an example, a spatially small feature with a huge variance in elevation compared to its surroundings would have a large persistence value, and not be simplified until high thresholds, whereas downsampling the data’s spatial resolution would quickly remove this feature.
Along with persistent homology, we utilize contour tree segmentation to create a hierarchical representation of the elevation data. We experimented with topological segmentations produced by the Morse-Smale complex (MSC) [18], merge, split, and contour trees, and found the contour tree to produce segments most closely aligned with flooded and dry regions while respecting elevation border pixels. Contour trees segment data by considering contours through a sweep of the function range, defining segments as the birth or split of a contour until it merges with another or vanishes. When simplifying by persistence, less persistent topological features are collapsed to the same elevation level as more persistent neighboring features. Because contour tree segments are regions bounded by contours at a given elevation level, this simplification corresponds to increasing the size of segments by collapsing the neighboring less persistent features into them. By considering multiple levels of simplification, this gives a convenient segmentation for selecting potential water levels in a flooded region, as water will naturally pool into depression features and stop at hill features or elevation contours.
@@
Similar topological techniques have previously been utilized for semi-automatic annotation in neuron tracing of large brain volumetric data [45]. This work utilizes the MSC to extract ridge-like structures which correspond to neuron center-lines, which can then guide user tracing. Simplification by persistence is used with a single empirically chosen persistence threshold to remove noise and create a sparser dataset for users to interact with. While taking inspiration from this work, our system adapts topological methods specifically to the labeling of flood extent in terrain data using contour tree segmentations. Furthermore, our system allows interactive selection of multiscale features by providing data simplified at varying persistence thresholds. This is necessary in this domain to find landforms at different scales within elevation data. Finally, FloodTrace is the only system the authors found that integrates topological segmentation into a web-based application providing unprecedented accessibility and functionalities to crowdworkers.
Figure 1: An illustration of the full processing pipeline for FloodTrace. Input is given to the application in corresponding RGB imagery and elevation data. The elevation data is used in the server backend to create a 3D mesh which is combined with the RGB imagery as shown in Figure 2. The elevation data is processed by the backend to create simplified topological segmentations at different levels of detail, as shown in Figure 3. The user then utilizes the elevation-guided tools described by subsection 3.2 on the rendering in the web frontend. These annotations can then be downloaded for training machine learning models directly in the case of expert-labeled data, or for aggregation in the case of crowdsourcing. Aggregated annotations can be given as input to our application to be used with our uncertainty visualization tool for review and correction, as shown in Figure 4, before being used to train machine learning models.